# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

### v3

@@ Appendix A Model Hyperparameters
Movielens 1M Netflix Prize GDSC1 CTRPv2 IMF DOTS Kiva MF Embedding Dimension 512 512 512 512 8 128 Learning Rate 0.001 0.001 0.001 0.001 0.001 0.0005 $L_{2}$ Weight Regularization 0.00001 0.00001 0.00001 0.00001 0.0001 0.0001 Early Stop Patience 10 10 10 10 10 10 Early Stop Delta 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 Bayesian SVD++ Embedding Dimension 32 - 10 10 10 10 MCMC Iterations 512 - 512 512 512 512 GC-MC Dropout 0.7 - 0.7 0.7 0.7 0.7 Hidden Layers 500, 75 - 500, 75 500, 75 500, 75 500, 75 Learning Rate 0.01 - 0.01 0.01 0.01 0.01 Training Epochs 3500 - 1000 1000 1000 1000 Mixture Model GCN Functions 2 - 2 2 2 2 GLocal-K Auto-Encoder Hiden Layers 500, 500 - 500, 500 500, 500 500, 500 500, 500 RBF Kernel Dimensions 5, 5 - 5, 5 5, 5 5, 5 5, 5 Global Kernel Dimension 3x3 - 3x3 3x3 3x3 3x3 $\lambda_{2}$ $L_{2}$ Weight Regularization 70 - 70 70 20 70 $\lambda_{s}$ $L_{2}$ Kernel Regularization 0.018 - 0.018 0.018 0.006 0.018 L-BFGS-B Pre-training Iterations 50 - 5 5 5 5 L-BFGS-B Fine-tuning Iterations 10 - 5 5 5 5
@@ Appendix B Additional experiments on Recommender Systems
This Appendix includes additional experiments (analysis of eccentricity-rating curves and evaluation of EAUC) in other popular Recommender Systems datasets: Movielens 100K and 10M [15], Douban Monti [34], and Tripadvisor [35]. The three former and the latter involve user-movie and client-restaurant rating prediction, respectively Table VI contains basic dataset information and statistics, while Table VII and Figure 9 showcase the eccentricity bias of ML models trained in each dataset. We hold the same conclusions as in results Section VI: a lower RMSE or MAE does not necessarily correlate with a lower eccentricity bias, and RMSE or MAE cannot reliably detect eccentricity bias in model performance; contrarily, both EAUC and its curve representation can accurately describe and measure it in all models and datasets and provide a bias-aware component of performance in dyadic regression tasks.
@@
Dataset Examples Users Items Density Movielens 100K 100,000 944 1,683 6.30% Movielens 10M 10,000,054 69,878 10,677 1.30% Douban Monti 136,066 3,000 3,000 1.50% Tripadvisor 491,170 25,388 12,384 0.20%
@@
Movielens 100K $D_{KS_{\mathcal{D}}}$: 0.415 Movielens 10M $D_{KS_{\mathcal{D}}}$: 0.437 RMSE MAE EAUC RMSE MAE EAUC Random 1.690 ± 0.009 1.381 ± 0.011 0.416 ± 0.015 1.841 ± 0.001 1.503 ± 0.000 0.409 ± 0.010 Dyad Average 0.978 ± 0.005 0.791 ± 0.005 0.401 ± 0.003 0.913 ± 0.001 0.719 ± 0.001 0.430 ± 0.007 Bayesian SVD++ 0.885 ± 0.005 0.693 ± 0.003 0.351 ± 0.006 0.759 ± 0.001 0.580 ± 0.001 0.328 ± 0.011 Glocal-K 0.882 ± 0.003 0.689 ± 0.004 0.343 ± 0.016 - - - GC-MC 0.891 ± 0.006 0.700 ± 0.005 0.371 ± 0.018 0.782 ± 0.001 0.594 ± 0.001 0.354 ± 0.013 MF 0.907 ± 0.008 0.711 ± 0.007 0.365 ± 0.022 0.782 ± 0.001 0.597 ± 0.001 0.341 ± 0.013
Douban Monti $D_{KS_{\mathcal{D}}}$: 0.466 Tripadvisor $D_{KS_{\mathcal{D}}}$: 0.539 RMSE MAE EAUC RMSE MAE EAUC Random 1.632 ± 0.011 1.333 ± 0.008 0.409 ± 0.019 1.826 ± 0.005 0.500 ± 0.005 0.425 ± 0.007 Dyad Average 0.771 ± 0.000 0.619 ± 0.000 0.416 ± 0.000 0.957 ± 0.004 0.737 ± 0.003 0.467 ± 0.020 Bayesian SVD++ 0.719 ± 0.009 0.564 ± 0.006 0.372 ± 0.004 0.942 ± 0.005 0.725 ± 0.004 0.453 ± 0.007 Glocal-K 0.723 ± 0.001 0.562 ± 0.001 0.382 ± 0.002 0.964 ± 0.015 0.742 ± 0.015 0.457 ± 0.049 GC-MC 0.732 ± 0.004 0.572 ± 0.006 0.399 ± 0.004 0.947 ± 0.012 0.733 ± 0.019 0.450 ± 0.029 MF 0.735 ± 0.002 0.574 ± 0.003 0.399 ± 0.001 0.972 ± 0.008 0.753 ± 0.007 0.448 ± 0.020

### v4

@@ Appendix A Model Hyperparameters
Movielens 1M Netflix Prize GDSC1 CTRPv2 IMF DOTS Kiva
MF Embedding Dimension 512 512 512 512 8 128
Learning Rate 0.001 0.001 0.001 0.001 0.001 0.0005
$L_{2}$ Weight Regularization 0.00001 0.00001 0.00001 0.00001 0.0001 0.0001
Early Stop Patience 10 10 10 10 10 10
Early Stop Delta 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
Bayesian SVD++ Embedding Dimension 32 - 10 10 10 10
MCMC Iterations 512 - 512 512 512 512
GC-MC Dropout 0.7 - 0.7 0.7 0.7 0.7
Hidden Layers 500, 75 - 500, 75 500, 75 500, 75 500, 75
Learning Rate 0.01 - 0.01 0.01 0.01 0.01
Training Epochs 3500 - 1000 1000 1000 1000
Mixture Model GCN Functions 2 - 2 2 2 2
GLocal-K Auto-Encoder Hiden Layers 500, 500 - 500, 500 500, 500 500, 500 500, 500
RBF Kernel Dimensions 5, 5 - 5, 5 5, 5 5, 5 5, 5
Global Kernel Dimension 3x3 - 3x3 3x3 3x3 3x3
$\lambda_{2}$ $L_{2}$ Weight Regularization 70 - 70 70 20 70
$\lambda_{s}$ $L_{2}$ Kernel Regularization 0.018 - 0.018 0.018 0.006 0.018
L-BFGS-B Pre-training Iterations 50 - 5 5 5 5
L-BFGS-B Fine-tuning Iterations 10 - 5 5 5 5
@@ Appendix B Additional experiments on Recommender Systems
This Appendix includes additional experiments (analysis of eccentricity-rating curves and evaluation of EAUC) in other popular Recommender Systems datasets: Movielens 100K and 10M [15], Douban Monti [39], and Tripadvisor [40]. The three former and the latter involve user-movie and client-restaurant rating prediction, respectively
Table VI contains basic dataset information and statistics, while Table VII and Figure 9 showcase the eccentricity bias of ML models trained in each dataset. We hold the same conclusions as in results Section VI: a lower RMSE or MAE does not necessarily correlate with a lower eccentricity bias, and RMSE or MAE cannot reliably detect eccentricity bias in model performance; contrarily, both EAUC and its curve representation can accurately describe and measure it in all models and datasets and provide a bias-aware component of performance in dyadic regression tasks.
@@
Dataset Examples Users Items Density
Movielens 100K 100,000 944 1,683 6.30%
Movielens 10M 10,000,054 69,878 10,677 1.30%
Douban Monti 136,066 3,000 3,000 1.50%
Tripadvisor 491,170 25,388 12,384 0.20%
@@
Movielens 100K $D_{KS_{\mathcal{D}}}$: 0.415 Movielens 10M $D_{KS_{\mathcal{D}}}$: 0.437
RMSE MAE EAUC RMSE MAE EAUC
Random 1.690 ± 0.009 1.381 ± 0.011 0.416 ± 0.015 1.841 ± 0.001 1.503 ± 0.000 0.409 ± 0.010
Dyad Average 0.978 ± 0.005 0.791 ± 0.005 0.401 ± 0.003 0.913 ± 0.001 0.719 ± 0.001 0.430 ± 0.007
Bayesian SVD++ 0.885 ± 0.005 0.693 ± 0.003 0.351 ± 0.006 0.759 ± 0.001 0.580 ± 0.001 0.328 ± 0.011
Glocal-K 0.882 ± 0.003 0.689 ± 0.004 0.343 ± 0.016 - - -
GC-MC 0.891 ± 0.006 0.700 ± 0.005 0.371 ± 0.018 0.782 ± 0.001 0.594 ± 0.001 0.354 ± 0.013
MF 0.907 ± 0.008 0.711 ± 0.007 0.365 ± 0.022 0.782 ± 0.001 0.597 ± 0.001 0.341 ± 0.013
Douban Monti $D_{KS_{\mathcal{D}}}$: 0.466 Tripadvisor $D_{KS_{\mathcal{D}}}$: 0.539
RMSE MAE EAUC RMSE MAE EAUC
Random 1.632 ± 0.011 1.333 ± 0.008 0.409 ± 0.019 1.826 ± 0.005 0.500 ± 0.005 0.425 ± 0.007
Dyad Average 0.771 ± 0.000 0.619 ± 0.000 0.416 ± 0.000 0.957 ± 0.004 0.737 ± 0.003 0.467 ± 0.020
Bayesian SVD++ 0.719 ± 0.009 0.564 ± 0.006 0.372 ± 0.004 0.942 ± 0.005 0.725 ± 0.004 0.453 ± 0.007
Glocal-K 0.723 ± 0.001 0.562 ± 0.001 0.382 ± 0.002 0.964 ± 0.015 0.742 ± 0.015 0.457 ± 0.049
GC-MC 0.732 ± 0.004 0.572 ± 0.006 0.399 ± 0.004 0.947 ± 0.012 0.733 ± 0.019 0.450 ± 0.029
MF 0.735 ± 0.002 0.574 ± 0.003 0.399 ± 0.001 0.972 ± 0.008 0.753 ± 0.007 0.448 ± 0.020